# Fast unbiased sampling of networks with given expected degrees and strengths

### v1

@@ 1 Introduction
The most widely used configuration model is the Chung-Lu model [4]. A critical limitation of the Chung-Lu model is that it oversamples edges between large-degree nodes [28, 17, 27]. While this bias is well-known and often considered negligible in sparse networks [5, 20], the bias is consequential in network analysis because large-degree nodes are a key determinant of network structure.
@@
We introduce a fast sampling algorithm for the configuration model for networks based on the maximum entropy principle. Building on the rejection sampling proposed by Miller and Hagberg (2011) [18], we adopt their method to work with the MaxEnt models by generating a set of candidate edges from an efficient proposal distribution, then selectively accepting these candidates to match the target probabilities. The resulting algorithm maintains statistical correctness while achieving dramatic computational speedup, making it applicable across undirected, directed, unweighted, and weighted enhanced configuration models. It should be noted that this paper focuses exclusively on the analysis and results for undirected networks, although the extension to the directed case is expected to be straightforward. Our algorithm makes theoretically rigorous configuration models practical for large network analysis, contributing to an accurate understanding of the structure induced by degree heterogeneity. A Python code for the proposed sampling algorithm can be found at https://github.com/EKUL-Skywalker/fast˙unbiased˙sampling˙of˙networks˙with˙given˙expected˙degrees˙and˙strengths.

### v3

@@ 1 Introduction
The most widely used configuration model is the Chung-Lu model [4]. A critical limitation of the Chung-Lu model is that it oversamples edges between large-degree nodes [29, 18, 28, 17]. While this bias is well-known and often considered negligible in sparse networks [5, 21], the bias is consequential in network analysis because large-degree nodes are a key determinant of network structure.
@@
We introduce a fast sampling algorithm for the configuration model for networks based on the maximum entropy principle. Building on the rejection sampling proposed by Miller and Hagberg (2011) [19], we adopt their method to work with the MaxEnt models by generating a set of candidate edges from an efficient proposal distribution, then selectively accepting these candidates to match the target probabilities. The resulting algorithm maintains statistical correctness while achieving dramatic computational speedup, making it applicable across undirected, directed, unweighted, and weighted enhanced configuration models. It should be noted that this paper focuses exclusively on the analysis and results for undirected networks, although the extension to the directed case is expected to be straightforward. Our algorithm makes theoretically rigorous configuration models practical for large network analysis, contributing to an accurate understanding of the structure induced by degree heterogeneity. A Python code for the proposed sampling algorithm can be found at https://github.com/EKUL-Skywalker/fastmaxent.
@@ 3.2 Weighted networks
4 Discussion
We have developed a fast sampling algorithm for the MaxEnt models by adapting the MH algorithm originally designed for the Chung-Lu model. Our evaluation demonstrated computational speedups of 10-100 fold in CPU time for unweighted networks and 10—100 fold for weighted networks relative to the brute-force method, while preserving statistical accuracy.
Our approach echoes the combinatorial adaptation of the MH algorithm [28] that replaces the underlying edge probability derived from a combinatorial analysis of edge occurrences between two nodes. We did not compare against the combinatorial MH algorithm, as we focus on sampling efficiency rather than model differences. We note that both fast algorithms make it computationally feasible to systematically compare different configuration models on large, degree-heterogeneous networks where such comparisons were previously prohibitive.
Several limitations remain. First, the inference cost for MaxEnt models is still ${\cal O}(N^{2})$. Second, we did not identify exact theoretical computational complexity of our sampling algorithm as it depends on the parameter distribution, i.e., the algorithm becomes most efficient when $\beta$ values are homogeneous, approaching $\beta_{\min}\simeq\beta_{i}+\beta_{j}$ for all node pairs $i,j$, as this ensures tight upper bounds throughout the sampling process.
Despite these limitations, our proposed sampling algorithm improved the efficiency of unbiased configuration models. We hope that our sampling algorithm will bridge the gap between theoretical rigor and computational practicality.